OpenStudy (anonymous):

how dp you simplify this problem? 3+1/4 / (5/6+2^2)

OpenStudy (phi):

is it \[\frac{3 +\frac{1}{4}}{\frac{5}{6}+2^2} \]?

OpenStudy (anonymous):

yes

OpenStudy (phi):

then do it step by step. first what is 2^2?

OpenStudy (anonymous):

4

OpenStudy (phi):

now what is 5/6 + 4 you first put the 4 over a common denominator of 6 like this \[ \frac{4}{1}\cdot \frac{6}{6}= \frac{24}{6}\] now add 5/6 + 24/6

OpenStudy (phi):

If you need more help with fractions, maybe this will help http://www.khanacademy.org/math/arithmetic/fractions/v/adding-and-subtracting-fractions

OpenStudy (anonymous):

sorry was writing it down at the same time

OpenStudy (anonymous):

29/6 is what i get

OpenStudy (phi):

OK, that is the bottom (denominator). Let's look at the top now. what is 3+1/4 ? (put 3 over a common denominator of 4 and then add)

OpenStudy (anonymous):

12/4

OpenStudy (anonymous):

add that with 1/4 and i get 13/4

OpenStudy (phi):

yes. now we use the great trick when dealing with fractions: if you multiply the top and bottom by the same thing, you do not change its value (that is how we changed 3 into 12/4 .... the whole number 3 can be thought of as having a 1 in the "bottom") so we have \[ \frac{\frac{13}{4}}{\frac{29}{6} }\] multiply top and bottom by 6/29

OpenStudy (anonymous):

do you find common or just mulitply

OpenStudy (anonymous):

where do you get that you multiply both by 6/29

OpenStudy (phi):

just multiply. like this \[ \frac{\frac{13}{4} \cdot \frac{6}{29}}{ \frac{29}{6} \cdot \frac{6}{29}}\]

OpenStudy (anonymous):

79/116 / 1 is what i get cause the bottom cancells out eash other right

OpenStudy (phi):

the reason we pick 6/29 is it makes the bottom equal 1, so we can ignore it anything divided by 1 is itself. so the problem becomes \[ \frac{13}{4} \cdot \frac{6}{29} \] Before multiplying it out I would divide the 4 by 2 and the 6 by 2 to simplify it \[ \frac{13}{2} \cdot \frac{3}{29} \]

OpenStudy (phi):

check your multiplication

OpenStudy (anonymous):

39/58

OpenStudy (phi):

yes, and I think that is in lowest (simplest) form.

OpenStudy (anonymous):

ok i was lost because when you go to freemathhelp.com they get 177/58

OpenStudy (phi):

If you multiplied out 13*6/ (4*29) you get 78/116 you can then simplify that to 39/58

OpenStudy (phi):

It depends exactly what the original problem is. Can you type in the problem with the correct parens, or show the original?

OpenStudy (anonymous):

give me a min i will try to copy and past the problem as it is wrote. i have to sign back in to get it out of the book.

OpenStudy (anonymous):

it wont let me but it is 3+1/4 for the num and 5/6+2^2 in the den

OpenStudy (anonymous):

I think that the way we did it is right because that is what i orig had for the answer

OpenStudy (phi):

I can get 177/58 if the original problem is \[ 3 + \frac{\frac{1}{4}}{\frac{5}{6}+2^2} \]

OpenStudy (anonymous):

how did you

OpenStudy (anonymous):

get that so i can put both down just in case

OpenStudy (phi):

If you use order of operations, that is how you would interpret what you posted: 3+1/4 / (5/6+2^2) however, so many people mean (3+1/4) / (5/6+2^2) that I assumed that was the problem.

OpenStudy (anonymous):

how did you get the other answer

OpenStudy (phi):

the bottom is the same as before= 29/6 so multiply top and bottom by 6/29 \[\frac{1}{4}\cdot \frac{6}{29} \]

OpenStudy (phi):

simplify by dividing the 6 and the 4 by 2, and you get 3/58

OpenStudy (phi):

now the problem is 3 + 3/58 put 3 over a common denominator of 58

OpenStudy (phi):

you get \[ \frac{174}{58}+\frac{3}{58} = \frac{177}{58} \]

OpenStudy (anonymous):

sorry got bumped i am going to look at what you put before i got bumped

OpenStudy (anonymous):

ty i finally got it do you have time to help with more

OpenStudy (anonymous):

x^2-2x+5x-(-x^2+7x)

OpenStudy (phi):

distribute the -1*(-x^2+7x) as the first step. that means multiply every term inside the parens by -1

OpenStudy (anonymous):

I get the neg and posit mixed up

OpenStudy (anonymous):

x^2+-7x i think

OpenStudy (phi):

-(-x^2+7x) is the same as -1*(-x^2+7x) which becomes -1* -x^2 + -1*7x - times - is + (2 minuses become a plus, a minus and a plus become a minus) so we get x^2 -7x

OpenStudy (phi):

or x^2 + -7x though people generally write it as x^2 -7x (fewer symbols) so now the problem is x^2-2x+5x+x^2 -7x combine "like terms" x^2 goes with x^2 , x with x,

OpenStudy (phi):

It sounds like it's tough out there, getting bumped and frozen!

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

looking over what you did

OpenStudy (anonymous):

so do you get get 2x^2+14x

OpenStudy (phi):

how did you get 14 x ? x^2-2x+5x+x^2 -7x we can "commute" the terms (change their order) x^2+x^2 + 5x -2x -7x if we think x^2 are cows and x are horses, how many cows and horses? or use this idea (x^2+x^2) + (5x -2x -7x) factor out x^2 from the first set of parens and x out of the second set (1+1)*x^2 + (5-2-7)*x

OpenStudy (anonymous):

not sure let me look

OpenStudy (anonymous):

the pos and neg confuse me is whyi get 2x^2+-4 i know that is not right i am still confused

OpenStudy (anonymous):

the 5-2-7 is what confuses me

OpenStudy (phi):

you mean 2x^2 + -4x (there is an x there, right?) or more simply 2x^2 -4x about the signs you can write , for example 8 + -4 or 8-4 they mean the same thing.

OpenStudy (phi):

5-2-7 means 5 minus 2 = 3 and then 3 minus 7 which is -4 or think of it as 5 + (-2) + (-7) you are adding 3 numbers

OpenStudy (phi):

here is a 5 minute video http://www.khanacademy.org/math/arithmetic/negative-numbers/v/adding-negative-numbers that might clear it up?

OpenStudy (anonymous):

ok that makes it easier so now we have 2x^2-4x

OpenStudy (phi):

yes. what to do next (if anything) depends on the question. often people would factor it to get 2x(x-2) but it depends...

OpenStudy (anonymous):

ok how did you get that

OpenStudy (anonymous):

it says to simplify

OpenStudy (anonymous):

i think what you got is what they want

OpenStudy (phi):

to factor 2x^2-4x notice that both terms can be divided by 2. so we divide both terms by 2, but put parens around the answer and put the 2 "out front" 2(x^2-2x) (we can "undo" this by multiplying each term by 2 to get 2x^2 -4 ) notice that both terms can be divided by x. Do the same thing 2x(x-2)

OpenStudy (anonymous):

do you have time for 1 more

OpenStudy (anonymous):

solve -9y-12(2-3y)=-25-2(y+6)

OpenStudy (phi):

*(we can "undo" this by multiplying each term by 2 to get 2x^2 -4x ) fixed a typo

OpenStudy (phi):

if you distribute the -12 on the left hand side, what do you get?

OpenStudy (anonymous):

-24+36y

OpenStudy (phi):

ok, now simplify the left side -9y-24+36y

OpenStudy (anonymous):

-24+27y

OpenStudy (phi):

ok, so we have, so far -24+27y = -25-2(y+6) distribute the -2 on the right. And write out the whole equation

OpenStudy (anonymous):

-25-2y-12 -24+27y=-25-2y-12

OpenStudy (phi):

I would combine the numbers on the right side: -25-12 = -37 add +24 to both sides of the equation. what do we get. Again write the whole equation.

OpenStudy (anonymous):

27y=-13-2y

OpenStudy (phi):

what do you think we should do next?

OpenStudy (anonymous):

add 2y to both sides

OpenStudy (phi):

we get?

OpenStudy (anonymous):

y=-13/29 i hope

OpenStudy (phi):

yes. looks good.

OpenStudy (anonymous):

ty soooo much i have more but dont want to stress you out lol

OpenStudy (phi):

post your questions, people will help you. but I am late as it is...

OpenStudy (anonymous):

ty again